Find all the errors in the following solution to an equation:
7/x = 8/x+3 +1
7(x+3) = 8x + 1
7x + 3 = 8x
-x = -3
x= 3

Answer by stanbon(75887)

Find all the errors in the following solution to an equation:
7/x = 8/x+3 +1
7(x+3) = 8x + 1
7x + 3 = 8x
-x = -3
x= 3
----------------
I'll do it the right way and let you find the errors:
7/x = 8/x+3 +1
Multiply thru by x(x+3) to get:
7(x+3) = 8x + x(x+3)
7x+21 = 8x + x^2 + 3x
7x+21 = 11x + x^2
x^2 + 4x -21 = 0
(x+7)(x-3) = 0
x = -7 or x = 3
=================
Cheers,
Stan H.